OpenStudy (anonymous):

A man has five pairs of socks (no two pairs are the same color). If he randomly selects two socks from a drawer, what is the probability that he gets a matching pair?

OpenStudy (anonymous):

1/2 chances or 5/10

OpenStudy (anonymous):

The answer in the book says 1/9...

OpenStudy (anonymous):

i never said it was gonna be right there was a half chance i was wrong

OpenStudy (anonymous):

I'm not trying to show that you're wrong... I just wanted to see if someone could explain this question to me. Because math is hard. And I have a test on probablility tomorrow.

OpenStudy (anonymous):

k there's 5 there it's an uneven number which means if u tryed to pair them all up 1 would be left behind

OpenStudy (anonymous):

lik if u got rid of that 1 sock and did 4x2 =8+1=9

OpenStudy (anonymous):

there's ur denominator

OpenStudy (anonymous):

he picks one sock it is some color

OpenStudy (anonymous):

then there are 9 socks left in the drawer, and only one has the same color as the sock he picked originally

OpenStudy (anonymous):

exactly it never said they were the same colors

OpenStudy (anonymous):

so the probability that the next sock he picks has the same color is \(\frac{1}{9}\)

OpenStudy (anonymous):

it is easier to think about it if you imagine picking one at a time instead of two at once

OpenStudy (anonymous):

we can also solve with thinking of picking two at once the number of ways he can pick the two socks out of the total of ten socks is \[\dbinom{10}{2}=\frac{10\times 9}{2}=5\times 9=45\] of those 45 ways, 5 of them are pairs that match, so you get the probability as \[\frac{5}{45}=\frac{1}{9}\]

OpenStudy (anonymous):

i think the first way is easier to envision, but it is a matter of taste i suppose